40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 19 00:07:16 2013, Jenne, Update, LSC, Green status 
    Reply  Tue Nov 19 01:11:19 2013, Jenne, Update, LSC, Green status 
    Reply  Tue Nov 19 14:47:59 2013, manasa, Update, LSC, Green status 
       Reply  Tue Nov 19 17:47:17 2013, Jenne, Update, LSC, PRMI+2arms attempt 
Message ID: 9413     Entry time: Tue Nov 19 17:47:17 2013     In reply to: 9411
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI+2arms attempt 

So far this afternoon, I have redone the IFO alignment, locked both arms with ALS, moved both arms off resonance, locked PRMI, and started bringing one arm back to resonance. 

The alignment was really not good, which I knew yesterday, but the ASS wasn't working yesterday.  I hand-did the alignment, and tried locking, which was easier with the slightly better alignment.

I locked both arms with ALS, found the resonances, and then moved them off resonance using Masayuki's scripts.

I then restored the PRM alignment, and locked the PRMI. 

I started bringing the Yarm back, but I kept losing lock when I got to about 0.1 transmission.


After losing lock several times, I switched over to looking at the ASS. I have figured out the problem, and fixed it.  The ASS for the arms now works again.

Looking at the StripTool plots of the lockin outputs for each arm, it was clear that the "L" traces were their usual size, but the "T" traces, which are demodulated versions of the transmission DC PDs, were tiny.  I investigated in the model, and the answer is obvious:  both the LSC and the ASS get the transmission information directly from the end sus computers.  Since we recently moved the normalization gain for the transmission diodes into the SUS models from the LSC model, this means that the ASS was seeing a differently sized signal than it had in the past. 

To fix this, I put a gain into the T_DEMOD_SIG filter banks for all 8 lockins that use info from the transmission DC PDs.  I used 1/g , where g is the gain that is in the C1:SUS-ETM#_TR#_GAIN channels.  For TRX, that number is -0.003, and for TRY that number is 0.002 .  So, in the .snap file that is used when turning on the ASS, I have given the Xarm lockins a gain of -333, and the Yarm lockins a gain of 500.  I chose this place, because the only thing that has happened to the signal until this point is a bandpass, so the rest of the servo gains can remain the same. 

I tested the ASS, and it works just like it used to.  I let it run, and align all of the optics, then I misaligned by a small amount each of the ETMs, saw that the lockin output values changed, and then were servoed back to zero.  So, it seems all good. 

ELOG V3.1.3-